# Covariant Measures of Non-Markovianity in Curved Spacetime

### v1

@@ Appendix C Wightman function with transverse velocity
Appendix D SDP formulation for the nearest Markovian comb
We briefly summarize the semidefinite program (SDP) used to compute the distance between a given process tensor $\Upsilon$ and the set $\mathcal{M}$ of Markovian (CP-divisible) combs, as discussed in Sec. III. We represent $\Upsilon$ as a positive operator on the multi-time input–output Hilbert space and measure its distance to $\mathcal{M}$ in trace norm (equivalently, in strategy norm up to normalization).
Let $\Upsilon_{M}\in\mathcal{M}$ denote a candidate Markovian comb. A standard SDP relaxation of the trace norm minimization problem
$\inf_{\Upsilon_{M}\in\mathcal{M}}\;\|\Upsilon-\Upsilon_{M}\|_{1}$ (62)
can be written as
minimize $\mathrm{Tr}[X]$
subject to $\begin{pmatrix}X&\Upsilon-\Upsilon_{M}\\ (\Upsilon-\Upsilon_{M})^{\dagger}&X\end{pmatrix}\succeq 0,$
$\Upsilon_{M}\in\mathcal{M},\qquad X\succeq 0,$ (63)
where $X$ is an auxiliary positive semidefinite variable. The block-positivity constraint in Eq. (63) is equivalent to $X\geq|\Upsilon-\Upsilon_{M}|$ and guarantees that the optimal value of $\mathrm{Tr}[X]$ coincides with $\|\Upsilon-\Upsilon_{M}\|_{1}$.
The Markovian set $\mathcal{M}$ is characterized either (i) by representing $\Upsilon_{M}$ as a product of single-step CPTP Choi operators and enforcing the causal normalization constraints step by step, or (ii) by imposing the linear constraints that implement CP-divisibility directly at the comb level, together with positivity and partial-trace identities (see Refs. [Sec. II] in the main text). Both approaches lead to a convex feasible set and are equivalent for the small system sizes considered here.
In practice, we implement the SDP using CVXPY (Python) or CVX (MATLAB) with a reliable interior-point solver such as MOSEK or SCS. For qubit detectors and moderate numbers of slots ($n\lesssim 4$55), the resulting SDPs are numerically tractable; the memory cost scales exponentially with $n$ because the Choi operator of an $n$-slot comb acts on a Hilbert space of dimension $d^{2n}$.
@@ Appendix D SDP formulation for the nearest Markovian comb
Appendix E Tensor-network mapping and convergence tests
For completeness we outline a nonperturbative numerical scheme that can be used to validate the perturbative memory-kernel analysis: a chain mapping of the scalar field followed by tensor-network simulation of the resulting open chain.
@@ Appendix E Tensor-network mapping and convergence tests
The field degrees of freedom coupled to the Unruh–DeWitt detector are mapped to a semi-infinite one-dimensional bosonic chain using standard orthogonal-polynomial techniques. After choosing a spectral density $J(\omega)$ and discretizing the continuum up to a cutoff frequency $\Lambda$, one performs a unitary transformation from the original mode basis to a set of effective chain modes with nearest-neighbour couplings. The resulting Hamiltonian takes the form
@@
where $H_{S}$ describes the detector, $H_{\text{chain}}$ is an approximately nearest-neighbour bosonic chain, and $H_{S\text{-chain}}$ couples the detector locally to the first site of the chain. Local bosonic Hilbert spaces are truncated to $d_{b}$ levels, and the discretization (frequency grid and orthogonal polynomials) is chosen to resolve both low- and high-frequency structure of $J(\omega)$ (e.g., logarithmic grids when appropriate).
@@
Time evolution of the detector+chain system is carried out using standard matrix-product-operator (MPO) or time-evolving block-decimation (TEBD) algorithms with bond dimension $\chi$. Convergence is assessed by increasing $\chi$ and the local cutoff $d_{b}$ until changes in relevant observables (e.g., the reduced detector state or the operational measure $N(\Upsilon)$ extracted from reconstructed kernels) fall below a preset threshold.
For the weak-coupling benchmarks considered here, typical parameter choices are
$\chi\in\{50,100,200\},\qquad d_{b}\in\{4,6,8\},\qquad\Delta t\sim 10^{-2},$ (65)
where $\Delta t$ is measured in units of the detector energy gap. The representative results quoted in the main text are obtained with $\chi=100$ and $d_{b}=6$ and have been spot-checked against $\chi=200$, $d_{b}=8$ for selected parameter sets, with discrepancies well below $10^{-3}$ in the quantities of interest. These tests indicate that, within the explored parameter regime, the perturbative kernel-based diagnostics and the nonperturbative tensor-network simulations are mutually consistent.

### v2

@@ Appendix C Wightman function with transverse velocity
Appendix D Semidefinite programs and numerical evaluation
Let $\Upsilon_{n:1}$ denote the Choi operator of an $n$-slot process tensor acting on the sequence of system input/output Hilbert spaces $\{\mathcal{H}^{\mathrm{i}}_{k},\mathcal{H}^{\mathrm{o}}_{k}\}_{k=1}^{n}$ (in the same slot ordering used throughout the main text). For any Hermitian operator
$\Delta\;:=\;\Upsilon_{n:1}-\widetilde{\Upsilon}_{n:1},$ (62)
the comb (strategy) norm $\|\Delta\|_{\mathrm{comb}}$ admits an operational interpretation as the optimal advantage in distinguishing the two processes using an arbitrary causal multi-time tester (co-strategy). Concretely, for two equiprobable hypotheses, the optimal success probability satisfies
$p_{\mathrm{succ}}^{\star}\;=\;\frac{1}{2}+\frac{1}{4}\,\|\Delta\|_{\mathrm{comb}}.$ (63)
Equivalently,
$\|\Delta\|_{\mathrm{comb}}\;=\;2\max_{\{T_{0},T_{1}\}}\Bigl\{\mathrm{Tr}\bigl[\Delta\,T_{0}\bigr]\Bigr\},$ (64)
where $\{T_{0},T_{1}\}$ is a two-outcome $n$-slot tester, and $T_{0},T_{1}$ are positive operators on the same tensor-product space as $\Upsilon_{n:1}$ satisfying the standard causal normalization constraints for testers.
A convenient SDP form uses the fact that $T:=T_{0}+T_{1}$ is a deterministic tester (co-strategy). Introducing intermediate normalization operators $T^{(k)}$ ($k=1,\dots,n$) and an initial state $\rho$ on $\mathcal{H}^{\mathrm{i}}_{1}$, the optimization (64) can be written as:
maximize: $\mathrm{Tr}\bigl[\Delta\,T_{0}\bigr]$ (65)
subject to: $T_{0}\succeq 0,\qquad T^{(n)}\succeq 0,\qquad 0\preceq T_{0}\preceq T^{(n)},$
$\mathrm{Tr}_{\mathrm{o}_{n}}\!\left[T^{(n)}\right]=\mathbb{I}_{\mathrm{i}_{n}}\otimes T^{(n-1)},$
$\mathrm{Tr}_{\mathrm{o}_{n-1}}\!\left[T^{(n-1)}\right]=\mathbb{I}_{\mathrm{i}_{n-1}}\otimes T^{(n-2)},$
$\hskip 42.00003pt\vdots$
$\mathrm{Tr}_{\mathrm{o}_{2}}\!\left[T^{(2)}\right]=\mathbb{I}_{\mathrm{i}_{2}}\otimes T^{(1)},$
$\mathrm{Tr}_{\mathrm{o}_{1}}\!\left[T^{(1)}\right]=\mathbb{I}_{\mathrm{i}_{1}}\otimes\rho,$
$\rho\succeq 0,\qquad\mathrm{Tr}[\rho]=1.$ (66)
With the normalization used in (63), the comb norm is then obtained as
[…]
We use (66) as the exact evaluation of $\|\cdot\|_{\mathrm{comb}}$ for fixed $\Delta$ in the small-$n$ regime where the SDP is tractable.
In practice, we implement the SDP using CVXPY (Python) or CVX (MATLAB) with a reliable interior-point solver such as MOSEK or SCS. For qubit detectors and moderate numbers of slots ($n\lesssim 4$-55), the resulting SDPs are numerically tractable; the memory cost scales exponentially with $n$ because the Choi operator of an $n$-slot comb acts on a Hilbert space of dimension $d^{2n}$.
@@ Appendix D Semidefinite programs and numerical evaluation
Appendix E Tensor-network mapping and convergence
For completeness we summarize a nonperturbative numerical scheme that can be used to benchmark the perturbative memory-kernel analysis: a chain mapping of the scalar field followed by tensor-network simulation of the resulting open chain.
@@ Appendix E Tensor-network mapping and convergence
The field degrees of freedom coupled to the Unruh–DeWitt detector are mapped to a semi-infinite one-dimensional bosonic chain using standard orthogonal-polynomial techniques. After specifying a spectral density $J(\omega)$ and discretizing the continuum up to a cutoff frequency $\Lambda$, one performs a unitary transformation from the original mode basis to a set of effective chain modes with nearest-neighbour couplings. The Hamiltonian then takes the form
@@
where $H_{S}$ describes the detector, $H_{\text{chain}}$ is an approximately nearest-neighbour bosonic chain, and $H_{S\text{-chain}}$ couples the detector locally to the first site of the chain. Local bosonic Hilbert spaces are truncated to $d_{b}$ levels, and the discretization (frequency grid and orthogonal polynomials) is chosen to resolve both low- and high-frequency features of $J(\omega)$ (e.g., logarithmic grids when appropriate).
@@
Time evolution of the detector+chain system is performed using standard matrix-product-operator (MPO) or time-evolving block-decimation (TEBD) algorithms with bond dimension $\chi$. Convergence is assessed by increasing $\chi$ and the local cutoff $d_{b}$ until changes in relevant observables (such as the reduced detector state or the operational measure $N(\Upsilon)$ extracted from reconstructed kernels) fall below a preset threshold.
For the weak-coupling benchmarks considered here, representative parameter choices are
$\chi\in\{50,100,200\},\quad d_{b}\in\{4,6,8\},\quad\Delta t\sim 10^{-2},$
with $\Delta t$ measured in units of the detector energy gap. The results quoted in the main text are obtained with $\chi=100$ and $d_{b}=6$, and have been spot-checked against $\chi=200$, $d_{b}=8$ for selected parameter sets, with discrepancies well below $10^{-3}$ in the quantities of interest. Within the explored regime, the perturbative kernel-based diagnostics and the nonperturbative tensor-network simulations are therefore mutually consistent.
Extension to stronger coupling.
Although the present work focuses on the weak-coupling regime, the tensor-network construction is, in principle, nonperturbative in $g$ and can be extended to explore stronger couplings $g\sim 1$, such as those relevant near horizons where large redshifts and enhanced backreaction may render the Born approximation unreliable. In practice, feasibility is controlled by the required chain length and bond dimension needed to capture longer correlation times, potentially necessitating advanced compression strategies or GPU-accelerated implementations. Breakdown of the second-order expansion is expected when higher-order terms in the Dyson series become comparable to the leading contribution, e.g., when $g^{2}\int\mathrm{d}\lambda\,\mathrm{d}\lambda^{\prime}\,C(\lambda,\lambda^{\prime})\sim 1,$ signalling the onset of significant non-Gaussian field effects and detector–field entanglement. In that regime, reconstructing $\Upsilon$ directly from the tensor-network evolution provides a natural nonperturbative continuation of the covariant framework developed here.